# LLM-Assisted Static Analysis for Detecting Security Vulnerabilities

### v1

@@ Abstract
Software is prone to security vulnerabilities. Program analysis tools to detect them have limited effectiveness in practice. While large language models (or LLMs) have shown impressive code generation capabilities, they cannot do complex reasoning over code to detect such vulnerabilities, especially because this task requires whole-repository analysis. In this work, we propose IRIS, the first approach that systematically combines LLMs with static analysis to perform whole-repository reasoning to detect security vulnerabilities. We curate a new dataset, CWE-Bench-Java, comprising 120 manually validated security vulnerabilities in real-world Java projects. These projects are complex, with an average of 300,000 lines of code and a maximum of up to 7 million. Out of 120 vulnerabilities in CWE-Bench-Java, IRIS detects 69 using GPT-4, while the state-of-the-art static analysis tool only detects 27. Further, IRIS also significantly reduces the number of false alarms (by more than 80% in the best case).
@@ 1 Introduction
Security vulnerabilities pose a major threat to the safety of software applications and its users. In 2023 alone, more than 29,000 CVEs were reported – almost 4000 higher than in 2022 [9]. Detecting vulnerabilities is extremely challenging despite advances in techniques to uncover them. A promising such technique called static taint analysis is widely used in popular tools such as GitHub CodeQL [1], Facebook Infer [11], Checker Framework [4], and Snyk Code [38]. These tools, however, face several challenges that greatly limit their effectiveness and accessibility in practice.
@@
False positives due to lack of precise context-sensitive and intuitive reasoning. Second, it is well-known that static analysis often suffers from low precision, i.e., it may often generate many false alarms [24, 21]. Such imprecision stems from multiple sources. For instance, the source or sink specifications may be spurious, or the analysis may over-approximate over branches in code or potential program inputs. Further, even if the specifications are correct, the context in which the detected source or sink is used may not lead to a vulnerability. Hence, in reality, a developer may need to triage through several (potentially false) security alerts, wasting developer time and effort.
Limitations of prior data-driven approaches to improve static taint analysis. Many techniques have been proposed to address the challenges of static taint analysis. For instance, Livshits et al. [33] proposed an automated approach, MERLIN, to mine specifications from library A more recent work, Seldon [6], improves the scalability of this approach by formulating the taint specification inference problem as a linear optimization task. However, such approaches rely on analyzing the code of third-party library to extract specifications, which is expensive and hard to scale. Researchers have also developed statistical and learning-based techniques to mitigate false positive alerts [23, 18, 35]. However, such approaches still have limited effectiveness in practice [24].
@@
Approach. We propose IRIS, the first neuro-symbolic approach for vulnerability detection that combines the strengths of static analysis and LLMs without suffering their limitations. Given a project to analyze for a given vulnerability class (or CWE), IRIS applies LLMs for mining CWE-specific taint specifications for third-party library APIs used in the project. IRIS augments such specifications with CodeQL,a static analysis tool, and uses it to detect security vulnerabilities. Our intuition here is because LLMs have seen numerous usages of such library APIs across projects, they have an understanding of which APIs are relevant for different CWEs.
Further, to address the imprecision problem of static analysis, we develop a contextual analysis technique with LLMs that drastically reduces the false positive alarms and minimizes the triaging effort for developers. Our key insight is that encoding the code-context and path-sensitive information in the prompt elicits more reliable reasoning from LLMs. Finally, our neuro-symbolic approach allows LLMs to do more precise whole-repository reasoning and minimizes the human effort involved in using static analysis tools. Figure 1 presents the end-to-end pipeline of IRIS.
Dataset. We curate a dataset of manually vetted and compilable Java projects, CWE-Bench-Java, containing 120 vulnerabilities (one per project) across four common vulnerability classes. The projects in the dataset are complex, containing 300K lines of code on average, and 10 projects with more than a million lines of code, making it a challenging benchmark for vulnerability detection.
Results. We evaluate IRIS on CWE-Bench-Java using eight diverse open- and closed-source LLMs. Overall, IRIS obtains the best results with GPT-4, detecting 69 vulnerabilities, which is 42 (35%) more than CodeQL. Among open-source LLMs, DeepSeekCoder 7B, despite being much smaller than other LLMs performs the best, detecting 67 vulnerabilities, followed by Llama 3 70B with 57 vulnerabilities. Further, our context-based filtering technique reduces false positive alerts by 80%.
@@ 2 Motivating Example
We illustrate the effectiveness of IRIS in detecting a previously known code-injection (CWE-094) vulnerability in cron-utils (ver. 9.1.5), a Java library for Cron data manipulation. Fig. 2(a) shows the relevant code snippets. A user-controlled string value passed into isValid function is transferred without sanitization to the parse function. If an exception is thrown, the function constructs an error message with the input. However, the error message is used to invoke method buildConstraintViolationWithTemplate of class ConstraintValidatorContext in javax.validator, which interprets the message string as a Java Expression Language (Java EL) expression. A malicious user may exploit this vulnerability by crafting a string containing a shell command such as Runtime.exec(‘rm -rf /’) to delete critical files on the server.
Detecting this vulnerability poses several challenges. First, the cron-utils library consists of 13K SLOC (lines of code excluding blanks and comments), which needs to be analyzed to find this vulnerability. This process requires analyzing data and control flow across several internal methods and third-party APIs. Second, the analysis needs to identify relevant sources and sinks. In this case, the value parameter of the public isValid method may contain arbitrary strings when invoked, and hence may be a source of malicious data. Additionally, external APIs like buildConstraintViolationWithTemplate can execute arbitrary Java EL expressions, hence they should be treated as sinks that are vulnerable to Code Injection attacks. Finally, the analysis also requires identifying any sanitizers that block the flow of untrusted data, as well as intuitive understanding of contextual information of the source and sink to suppress false positives.
@@
In contrast, IRIS takes a different approach by inferring project- and vulnerability-specific specifications on-the-fly by using LLMs. As shown in Fig. 2(b), with simple prompts, the LLM-based components in IRIS correctly identify the untrusted source and the vulnerable sink. IRIS augments CodeQL with these specifications and successfully detects the unsanitized data-flow path between the detected source and sink in the repository. However, in addition to the true vulnerable path, augmented CodeQL produces many false positives, which are hard to eliminate using concrete rules. To solve this challenge, IRIS encodes the detected code paths and the surrounding context into a simple prompt and uses an LLM to classify it as true or false positive. Specifically, out of 8 paths reported by static analysis, 5 false positives are filtered out, leaving the path in Fig. 2(a) as one of the final alarms. While we use GPT 4 for the above example, we show that one can easily adapt a wide range of LLMs for effective vulnerability detection by using zero- and few-shot prompting techniques. Overall, we observe that IRIS can detect many such vulnerabilities that are beyond the reach of CodeQL-like static analysis tools, while keeping false alarms to a minimum.
@@ 3.2 Candidate Source/Sink Extraction
A project may use various third-party APIs whose specifications (i.e., source or sink) may be unknown – reducing the effectiveness of taint analysis. In addition, internal APIs might accept untrusted input from downstream libraries as well. Hence, our goal is to automatically infer specifications for such APIs. We define a specification as a 3-tuple , where is the type of node to match in , is an N-tuple of strings describing the package, class, method name, signature, and argument/parameter position (if applicable) of an API, and is the role of the API. For example, the specification denotes that the first argument of exec method of Runtime class is a sink for a vulnerability class (OS command injection). A static analysis tool maps these specifications to sets of nodes or in .
To identify taint specifications and , we first extract : external (third-party) library APIs that are invoked in the given Java project and are potential candidates to be taint sources or sinks. We also extract , internal library APIs that are public and may be invoked by a downstream library. We utilize CodeQL’s query language to extract such candidates and their corresponding metadata such as method name, type signature, enclosing packages and classes, and even JavaDoc documentations, if applicable.
@@ 3.3 Inferring Taint Specifications using LLMs
We develop an automated specification inference technique: , where are candidate specifications for sources and sinks. In this work, we do not consider sanitizer specifications, because they typically do not vary for the vulnerability classes that we consider. We use LLMs to infer taint specifications. Specifically, external APIs in can be classified as either source or sink, while internal APIs in can have their formal parameters identified as sources. The left and middle panel of Fig. 2(b) show partial user prompts for inferring source and sink specifications from external APIs and internal function formal parameters.
@@ 3.4 Vulnerability Detection
Once we obtain all the source and sink specifications from the LLM, the next step is to combine it with a static analysis engine to detect vulnerable paths, i.e., , in a given project. In this work, we use CodeQL [14] for this step. CodeQL represents programs as data flow graphs and provides a query language, akin to Datalog [37], to analyze such graphs. Many security vulnerabilities and bugs can be modeled using queries written in CodeQL and can be executed against data flow graphs extracted from such programs. Given a data flow graph of a project , CWE-specific source and sink specifications, and a query for a given vulnerability class , CodeQL returns a set of unsanitized paths in the program. Formally,
@@ 3.5 Triaging of Alerts via Contextual Analysis
Figure 4 presents an example prompt for contextual analysis. The prompt includes CWE information and code snippets for nodes along the path, with an emphasis on the source and sink. Specifically, we include lines surrounding the exact source and sink location, as well as the enclosing function and class. The exact line of source and sink is marked with a comment. For the intermediate steps, we include the file names and the line of code. When the path is too long, we keep only a subset of nodes to limit the size of the prompt. As such, we provide the full context for the potential vulnerability to be thoroughly analyzed.
@@ 3.6 Evaluation Metrics
We evaluate the performance of IRIS using one key metrics, VulDetected. For evaluation, we assume that the label for a project is provided as a set of crucial program points where the vulnerable paths should pass through. In practice, these are typically the patched methods that can be collected from each vulnerability report. If at least one detected vulnerable path passes through a fixed location for the given vulnerability, then we consider the vulnerability detected. Suppose is the set of detected and/or filtered paths from prior stages, the metrics, including #Vuls the total number of detected vulnerabilities across a set of projects, is formally defined as:
@@ 4 CWE-Bench-Java: A Dataset of Security Vulnerabilities in Java
As shown in the statistics (Fig. 5), the sheer size of these projects make them challenging to analyze for any static analysis tool or ML-based tool. Each project in CWE-Bench-Java comes with GitHub information, vulnerable and fix version, CVE metadata, a script that automatically fetches and builds, and the set of program locations that involve the vulnerability. Our tools are easily extensible and can be directly used to extract vulnerabilities with other CWEs.
CWE | #Vuls | Detected | ||||||||
QL | IRIS with | |||||||||
G 7B | L3 7B | Mi 7B | DS 7B | DS 33B | L3 70B | GPT 3.5 | GPT 4 | |||
22 | 55 | 22 | 3 () | 19 () | 28 () | 28 () | 10 () | 29 () | 27 () | 34 () |
78 | 13 | 1 | 0 () | 3 () | 4 () | 6 () | 8 () | 2 () | 1 () | 7 () |
79 | 31 | 4 | 11 () | 13 () | 11 () | 19 () | 20 () | 16 () | 15 () | 17 () |
94 | 21 | 0 | 8 () | 12 () | 6 () | 14 () | 7 () | 10 () | 10 () | 11 () |
Total | 120 | 27 | 22 () | 47 () | 49 () | 67 () | 45 () | 57 () | 53 () | 69 () |
@@ 5 Evaluation
RQ 2: How effective is the contextual filtering approach in reducing false positive alerts?
@@ 5.1 Experimental Setup
LLM selection and CodeQL baseline. We select two closed-source LLMs from OpenAI: GPT 4 (gpt-4-0125-preview) and GPT 3.5 (gpt-3.5-turbo-0125) for our evaluation. We also select instruction-tuned versions of six state-of-the-art open-source LLMs via huggingface API: Llama 3 8B and 70B, DeepSeekCoder 7B and 33B, Mistral 7B, and Gemma 7B. For baseline, we use CodeQL 2.15.3 and its built-in Security queries specifically designed for each CWE.
5.2 RQ1: Vulnerability Detection Effectiveness of IRIS
The results, as summarized in Table 1, demonstrate the superior performance of IRIS across all CWEs with most LLMs, compared to CodeQL. Specifically, IRIS successfully identifies 69 vulnerabilities with GPT-4, 42 more than CodeQL. While GPT-4 excels the most, smaller but more specialized LLM like DeepSeekCoder 8B can also detect 67. To further contextualize the significance of results, we note that the dataset contains many exceptionally large and complex codebases such as Apache Camel, which contains over 1M SLOC. Notably with GPT-4, IRIS identifies the vulnerability within Camel, further underscoring its scalability. The primary reason for such gain in performance is due to the relevant taint specifications that are inferred by LLMs. Except for Gemma 7B, we observe that IRIS obtains significant better results than CodeQL with all other LLMs.
Effectiveness across LLMs. Out of 120 vulnerabilities, 34 are not detected by IRIS with any LLM, 78 are detected by at least two LLMs, and 10 are detected by all. In fact, all but one vulnerability detected by CodeQL cannot be detected by IRIS with GPT-4. The undetected case is due to a missing source specification related to Java annotations, which we currently do not support. Moreover, 8 vulnerabilities are detected by only one LLM: GPT-4 (1), Mistral 7B (1), DeepSeekCoder 7B (4), and DeepSeekCoder 33B (2). Observing only small discrepancies, we conclude that LLMs share a lot of common knowledge and are generally applicable to specification inference.
Effectiveness across CWEs. We observe that CWE-78 (OS Command Injection) is particularly challenging to detect for most LLMs. Our manual investigation revealed that the vulnerability patterns in CWE-78 are particularly complex, involving injection of OS commands through gadget-chains [2] or external side-effects like file writes, which cannot be tracked using static analysis. This underscores the fundamental limitations of static analysis (vs dynamic), which is beyond the scope of this work. Notwithstanding, we find IRIS to improve recall across all CWEs.
5.3 RQ2: Effectiveness of Contextual Analysis Approach in IRIS
We next study the effectiveness of contextual analysis in reducing false alarms. While filtering paths with contextual analysis can reduce false alarms, it may also potentially reduce the number of detected vulnerabilities. Table 2 presents the average number of reported paths per project and the number of detected vulnerabilities that remain after filtering. Due to the large number of paths per project, we only perform this experiment on a randomly selected set of 40 projects (10 per CWE) across the three best LLMs from Section 5.2. In all cases, we see a minor drop in #Vuls but a much more significant drop in average #Paths reported, up to a reduction rate of 81% with GPT-4.
Better reasoning capability in larger LLMs helps contextual analysis. Our manual investigation reveals that GPT-4 is very good at reasoning about the predicted paths, which aligns with its highest reduction rate in Table 2. GPT-4 not only reasons about the trustworthiness of taint sources but also understands the developer’s intent using natural language information in code and comments, which allows it to filter out many false positives. We also observe that when true positives are marked false by GPT-4, it is typically because the given context is insufficient for analysis. However, based on the path information alone, GPT-4 can already perform a very precise and reasonable analysis. It remains to be seen how further context improves precision, but we leave it for future work.
CWE | CodeQL | DeepSeekCoder 7B | Llama 3 70B | GPT-4 | ||||
#Paths | #Vuls | #Paths (%Red.) | #Vuls | #Paths (%Red.) | #Vuls | #Paths (%Red.) | #Vuls | |
22 | 62 | 6 | 282 ( %) | 7 ( ) | 152 ( %) | 7 ( ) | 93 ( %) | 10 ( ) |
78 | 4 | 1 | 347 ( %) | 4 ( ) | 24 ( %) | 2 ( ) | 22 ( %) | 3 ( ) |
79 | 32 | 2 | 1002 ( %) | 10 ( ) | 406 ( %) | 10 ( ) | 350 ( %) | 10 ( ) |
94 | 0 | 0 | 1037 ( %) | 7 ( ) | 795 ( %) | 7 ( ) | 270 ( %) | 6 ( ) |
All | 26 | 9 | 648 ( %) | 28 ( ) | 328 ( %) | 26 ( ) | 179 ( %) | 29 ( ) |
5.4 RQ3: Quality of Inferred Source and Sink Specifications
CWE | #Cand. | GPT-4 | DS 7B | ||
%S | %N | %S | %N | ||
22 | 130,974 | 2.03% | 1.90% | 4.27% | 4.01% |
78 | 25,605 | 4.73% | 3.37% | 3.67% | 3.33% |
79 | 37,138 | 5.69% | 4.69% | 4.28% | 4.56% |
94 | 36,325 | 5.12% | 7.83% | 6.11% | 6.21% |
Total | 230,042 | 3.41% | 3.45% | 4.50% | 4.37% |
Next, we analyze whether LLMs are capable of inferring high-quality source and sink specifications. Table 3 presents the statistics on LLM-inferred specifications, showing that about 4% of candidates are labeled on average. We perform manual analysis on ten random samples each of source and sink specifications, per combination of CWE and LLM, and estimate the overall precision of specifications (Fig. 6). Clearly, GPT-4 obtains significantly higher precision (over 70%) than other LLMs.
Over-approximate specifications are beneficial. While the precision for other LLMs is lower, we note that the over-approximation is, in fact, a partial solution to overcome a fundamental limitation of CodeQL, viz., limited taint specifications. LLMs thus have the potential to expand the coverage of taint analysis. However, we also note that the effect of this imprecision can be mitigated by a lot via contextual analysis, which we presented in Sec. 5.3.
Continuous taint specification inference is necessary. Our results show that there is a high number of both unique and recurring sources and sinks: about 900 per CWE unique and 700 recurring (tables in Appendix). This indicates that even if previously inferred specifications are useful, a significant number of new relevant APIs still remain and need to be labeled for effective vulnerability detection. This observation strongly motivates the design of IRIS that infers these specifications on-the-fly for each project via LLMs, instead of relying on a fixed corpus of specifications like CodeQL.

### v2

@@ Abstract
Software is prone to security vulnerabilities. Program analysis tools to detect them have limited effectiveness in practice due to their reliance on human labeled specifications. Large language models (or LLMs) have shown impressive code generation capabilities but they cannot do complex reasoning over code to detect such vulnerabilities especially since this task requires whole-repository analysis. We propose IRIS, a neuro-symbolic approach that systematically combines LLMs with static analysis to perform whole-repository reasoning for security vulnerability detection. Specifically, IRIS leverages LLMs to infer taint specifications and perform contextual analysis, alleviating needs for human specifications and inspection. For evaluation, we curate a new dataset, CWE-Bench-Java, comprising 120 manually validated security vulnerabilities in real-world Java projects. A state-of-the-art static analysis tool CodeQL detects only 27 of these vulnerabilities whereas IRIS with GPT-4 detects 55 () and improves upon CodeQL’s average false discovery rate by 5% points. Furthermore, IRIS identifies 6 previously unknown vulnerabilities which cannot be found by existing tools.
@@ 1 Introduction
Security vulnerabilities pose a major threat to the safety of software applications and its users. In 2023 alone, more than 29,000 CVEs were reported—almost 4000 higher than in 2022 [9]. Detecting vulnerabilities is extremely challenging despite advances in techniques to uncover them. A promising such technique called static taint analysis is widely used in popular tools such as GitHub CodeQL [1], Facebook Infer [11], Checker Framework [4], and Snyk Code [39]. These tools, however, face several challenges that greatly limit their effectiveness and accessibility in practice.
@@
False positives due to lack of precise context-sensitive and intuitive reasoning. Second, it is well-known that static analysis often suffers from low precision, i.e., it may generate many false alarms [24, 21]. Such imprecision stems from multiple sources. For instance, the source or sink specifications may be spurious, or the analysis may over-approximate over branches in code or possible inputs. Further, even if the specifications are correct, the context in which the detected source or sink is used may not be exploitable. Hence, a developer may need to triage through several potentially false security alerts, wasting significant time and effort.
Limitations of prior data-driven approaches to improve static taint analysis. Many techniques have been proposed to address the challenges of static taint analysis. For instance, [34] proposed a probabilistic approach, MERLIN, to automatically mine taint specifications. A more recent work, Seldon [6], improves the scalability of this approach by formulating the taint specification inference problem as a linear optimization task. However, such approaches rely on analyzing the code of third-party libraries to extract specifications, which is expensive and hard to scale. Researchers have also developed statistical and learning-based techniques to mitigate false positive alerts [23, 18, 36]. However, such approaches still have limited effectiveness in practice [24].
@@
Approach. We propose IRIS, a neuro-symbolic approach for vulnerability detection that combines the strengths of static analysis and LLMs. Fig. 1 presents an overview of IRIS. Given a project to analyze for a given vulnerability class (or CWE), IRIS applies LLMs for mining CWE-specific taint specifications. IRIS augments such specifications with CodeQL, a tool for static taint analysis. Our intuition here is because LLMs have seen numerous usages of such library APIs, they have an understanding of the relevant APIs for different CWEs. Further, to address the imprecision problem of static analysis, we propose a contextual analysis technique with LLMs that reduces the false positive alarms and minimizes the triaging effort for developers. Our key insight is that encoding the code-context and path-sensitive information in the prompt elicits more reliable reasoning from LLMs. Finally, our neuro-symbolic approach allows LLMs to do more precise whole-repository reasoning and minimizes the human effort involved in using static analysis tools.
Dataset. We curate a dataset of manually vetted and compilable Java projects, CWE-Bench-Java, containing 120 vulnerabilities (one per project) across four common vulnerability classes. The projects in the dataset are complex, containing 300K lines of code on average, and 10 projects with more than a million lines of code, making it a challenging benchmark for vulnerability detection. Our code and dataset are in the supplementary material and will be open-sourced upon publication.
Results. We evaluate IRIS on CWE-Bench-Java using 8 diverse open- and closed-source LLMs. Overall, IRIS obtains the best results with GPT-4, detecting 55 vulnerabilities, which is 28 () more than CodeQL, the existing best-performing static analyzer. We show that the increase is not at the expense of false positives, as IRIS with GPT-4 achieves an average false discovery rate of , which is lower than that of CodeQL. Further, when applied to the latest versions of 30 Java projects, IRIS with GPT-4 discovered 6 previously unknown vulnerabilities.
@@ 2 Motivating Example
We illustrate the effectiveness of IRIS in detecting a previously known code-injection (CWE-094) vulnerability in cron-utils (ver. 9.1.5), a Java library for Cron data manipulation. Fig. 2 shows the relevant code snippets. A user-controlled string value passed into isValid function is transferred without sanitization to the parse function. If an exception is thrown, the function constructs an error message with the input. However, the error message is used to invoke method buildConstraintViolationWithTemplate of class ConstraintValidatorContext in javax.validator, which interprets the message string as a Java Expression Language (Java EL) expression. A malicious user may exploit this vulnerability by crafting a string containing a shell command such as Runtime.exec(‘rm -rf /’) to delete critical files on the server.
Detecting this vulnerability poses several challenges. First, the cron-utils library consists of 13K SLOC (lines of code excluding blanks and comments), which needs to be analyzed to find this vulnerability. This process requires analyzing data and control flow across several internal methods and third-party APIs. Second, the analysis needs to identify relevant sources and sinks. In this case, the value parameter of the public isValid method may contain arbitrary strings when invoked, and hence may be a source of malicious data. Additionally, external APIs like buildConstraintViolationWithTemplate can execute arbitrary Java EL expressions, hence they should be treated as sinks that are vulnerable to Code Injection attacks. Finally, the analysis also requires identifying any sanitizers that block the flow of untrusted data.
@@
In contrast, IRIS takes a different approach by inferring project- and vulnerability-specific specifications on-the-fly by using LLMs. The LLM-based components in IRIS correctly identify the untrusted source and the vulnerable sink. IRIS augments CodeQL with these specifications and successfully detects the unsanitized data-flow path between the detected source and sink in the repository. However, augmented CodeQL produces many false positives, which are hard to eliminate using logical rules. To solve this challenge, IRIS encodes the detected code paths and the surrounding context into a simple prompt and uses an LLM to classify it as true or false positive. Specifically, out of 8 paths reported by static analysis, 5 false positives are filtered out, leaving the path in Fig. 2 as one of the final alarms. Overall, we observe that IRIS can detect many such vulnerabilities that are beyond the reach of CodeQL-like static analysis tools, while keeping false alarms to a minimum.
@@ 3.2 Candidate Source/Sink Extraction
A project may use various third-party APIs whose specifications may be unknown—reducing the effectiveness of taint analysis. In addition, internal APIs might accept untrusted input from downstream libraries. Hence, our goal is to automatically infer specifications for such APIs. We define a specification as a 3-tuple , where is the type of node to match in , is an N-tuple of strings describing the package, class, method name, signature, and argument/parameter position (if applicable) of an API, and is the role of the API. For example, the specification denotes that the first argument of exec method of Runtime class is a sink for a vulnerability class (OS command injection). A static analysis tool maps these specifications to sets of nodes or in .
To identify taint specifications and , we first extract : external library APIs that are invoked in the given Java project and are potential candidates to be taint sources or sinks. We also extract , internal library APIs that are public and may be invoked by a downstream library. We use CodeQL to extract such candidates and their corresponding metadata such as method name, type signature, enclosing packages and classes, and even JavaDoc documentations, if applicable.
@@ 3.3 Inferring Taint Specifications using LLMs
We develop an automated specification inference technique: , where are candidate specifications for sources and sinks. In this work, we do not consider sanitizer specifications, because they typically do not vary for the vulnerability classes that we consider. We use LLMs to infer taint specifications. Specifically, external APIs in can be classified as either source or sink, while internal APIs in can have their formal parameters identified as sources. In the Appendix, we show the user prompts for inferring source and sink specifications from external APIs and internal function formal parameters.
@@ 3.4 Vulnerability Detection
Once we obtain all the source and sink specifications from the LLM, the next step is to combine it with a static analysis engine to detect vulnerable paths, i.e., , in a given project. In this work, we use CodeQL [14] for this step. CodeQL represents programs as data flow graphs and provides a query language, akin to Datalog [38], to analyze such graphs. Many security vulnerabilities can be modeled using queries written in CodeQL and can be executed against data flow graphs extracted from such programs. Given a data flow graph of a project , CWE-specific source and sink specifications, and a query for a given vulnerability class , CodeQL returns a set of unsanitized paths in the program. Formally,
@@ 3.5 Triaging of Alerts via Contextual Analysis
Fig. 4 presents an example prompt for contextual analysis. The prompt includes CWE information and code snippets for nodes along the path, with an emphasis on the source and sink. Specifically, we include lines surrounding the exact source and sink location, as well as the enclosing function and class. The exact line of source and sink is marked with a comment. For the intermediate steps, we include the file names and the line of code. When the path is too long, we keep only a subset of nodes to limit the size of the prompt. As such, we provide the full context for the potential vulnerability to be thoroughly analyzed.
@@ 3.6 Evaluation Metrics
We evaluate the performance of IRIS and its baselines using three key metrics: number of vulnerability detected (#Detected), average false discovery rate (AvgFDR), and average F1 (AvgF1). For evaluation, we assume that we have a dataset where each is a Java project, and known to contain at least one vulnerability. The label for a project is provided as a set of crucial program points where the vulnerable paths should pass through, indicated by . In practice, these are typically the patched methods that can be collected from each vulnerability report. If at least one detected vulnerable path passes through a fixed location for the given vulnerability, then we consider the vulnerability detected. Let be the set of detected paths for each project from prior stages. The metrics are formally defined as follows:
We note that might sometimes be undefined due to division-by-zero if the detection tool retrieves no path (). Therefore, for AvgFDR to be meaningful, we only consider the projects where the detection tool produces at least one positive result. On the other hand, AvgF1 would not suffer from this problem because when no positive result is returned, making the whole F1 term regardless of undefinedness of .
@@ 4 CWE-Bench-Java: A Dataset of Security Vulnerabilities in Java
As shown in the statistics (Fig. 5), the sheer size of these projects make them challenging to analyze for any static analysis tool or ML-based tool. Each project in CWE-Bench-Java comes with GitHub information, vulnerable and fix version, CVE metadata, a script that automatically fetches and builds, and the set of program locations that involve the vulnerability.
@@ 5 Evaluation
RQ 2: Does IRIS detect new, previously unknown vulnerabilities?
@@ 5.1 Experimental Setup
We select two closed-source LLMs from OpenAI: GPT 4 (gpt-4-0125-preview) and GPT 3.5 (gpt-3.5-turbo-0125) for our evaluation. We also select instruction-tuned versions of three open-source LLMs via huggingface API: Llama 3 8B and 70B, and DeepSeekCoder 7B. For the CodeQL baseline, we use version 2.15.3 and its built-in Security queries specifically designed for each CWE. Other baselines included are Facebook Infer [11], SpotBugs [26], and Snyk [39]. We expand further on the other experimental setups in the appendix.
5.2 RQ1: Effectiveness of IRIS on Detecting Existing Vulnerabilities
Method | #Detected (/120) | Detection Rate () | Avg FDR () | Avg F1 Score | |
CodeQL | |||||
IRIS + | GPT-4 | () | () | () | () |
GPT-3.5 | () | () | () | () | |
L3 8B | () | () | () | () | |
L3 70B | () | () | () | () | |
DSC 7B | () | () | () | () |
Effectiveness of IRIS. The results in Table 1 highlight IRIS’s superior performance compared to CodeQL. Specifically, IRIS, when paired with GPT-4, identifies 55 vulnerabilities—28 more than CodeQL. While GPT-4 shows the highest efficacy, smaller, specialized LLMs like DeepSeekCoder 7B still detect 52 vulnerabilities, suggesting that our approach can effectively leverage smaller-scale models, enhancing accessibility. Notably, this increase in detected vulnerabilities does not compromise precision, as evidenced by IRIS’s lower average false discovery rate (FDR) with GPT-4 compared to CodeQL. Moreover, IRIS improves average F1 by 0.1, reflecting a better balance between precision and recall. We note that the reported average FDR is an upper bound, as our metrics may overlook other true vulnerabilities in the repository. To further assess detection accuracy, we randomly sampled 50 alarms reported by IRIS using GPT-4, and found that 27 out of 50 exhibit potential attack surfaces, yielding a more refined estimated false discovery rate of 46%.
Table 2 presents a detailed breakdown of detected vulnerabilities, comparing IRIS against various baselines. With the exception of IRIS using Llama-3 8B, which underperforms in detecting CWE-22 vulnerabilities, IRIS consistently outperforms all other baselines. Notably, CWE-78 (OS Command Injection) remains particularly challenging for all LLMs. Our manual investigation revealed that the vulnerability patterns in CWE-78 are highly intricate, often involving OS command injections via gadget-chains [2] or external side effects, such as file writes, which are difficult to track using static analysis. This highlights the inherent limitations of static analysis, as opposed to dynamic approaches—an area that we leave for future work.
CWE | #Vuls | Baselines | IRIS with | |||||||
QL | Infer | SB | Snyk | GPT-4 | GPT-3.5 | L3 8B | L3 70B | DSC 7B | ||
CWE-22 | 55 | 22 | 0 | 2 | 21 | 31 () | 25 () | 19 () | 29 () | 25 () |
CWE-78 | 13 | 1 | 0 | 1 | 1 | 3 () | 1 () | 2 () | 2 () | 3 () |
CWE-79 | 31 | 4 | 0 | 1 | 1 | 13 () | 13 () | 9 () | 14 () | 14 () |
CWE-94 | 21 | 0 | 0 | 0 | 0 | 8 () | 8 () | 11 () | 9 () | 10 () |
All | 120 | 27 | 0 | 4 | 23 | 55 () | 47 () | 41 () | 54 () | 52 () |
5.3 RQ2: Previously Unknown Vulnerabilities by IRIS
We applied IRIS with GPT-4 to the latest versions of 30 Java projects. Among the 16 inspected projects where IRIS raised at least one alert, we identified 6 potential vulnerabilities, of which 4 have been reported to the developers and are pending confirmation. These reported vulnerabilities include 3 instances of path injection (CWE-22) and one case of cross-site scripting (CWE-94). To ensure that these vulnerabilities were indeed uncovered due to IRIS’s integration with LLMs, we verified that they were not detectable by CodeQL alone. Detailed findings are presented in the appendix, but we highlight one such vulnerability in Fig. 8. CodeQL was unable to detect this issue due to a missing source specification, while GPT-4 successfully flagged the API endpoint restoreFromCheckpoint as a potential entry point for attack.
5.4 RQ3: Quality of LLM-Inferred Taint Specifications
The LLM-inferred taint specifications are fundamental to IRIS’s effectiveness. To assess the quality of these specifications, we conducted two experiments. First, we used CodeQL’s taint specifications as a benchmark to estimate the recall of both source and sink specifications inferred by LLMs (Fig. 6). However, since CodeQL offers a limited set of specifications, we also needed to assess the quality of inferred specifications outside of its known coverage. To this end, we manually analyzed 960 randomly selected samples of LLM-inferred source and sink labels (30 per combination of CWE and LLM) and estimated the overall precision of the specifications (Fig. 7).
LLM-inferred sinks can replace CodeQL sinks. Overall, LLMs demonstrated high recall when tested against CodeQL’s sink specifications (Fig. 6), with GPT-4 scoring the highest (). While the recall for source specifications was generally lower, we found that CodeQL tends to over-approximate its source specifications to compensate for a low detection rate. On the other hand, GPT-4 achieved high precision (over 70%) in manual evaluations (Fig. 7), aligning with the lower false discovery rate previously reported in Table 1. For other LLMs, the combination of high recall but lower precision suggests a tendency to over-approximate sink specifications.
Over-approximating specifications can benefit IRIS. Although the precision for LLMs other than GPT-4 is lower, over-approximation can actually help address a core limitation of CodeQL—its restricted set of taint specifications. By over-approximating, LLMs expand the coverage of taint analysis, offering a partial solution to CodeQL’s limited scope. The impact of this imprecision can be mitigated through contextual analysis as we show next in the ablation studies.
5.5 RQ4: Ablation Studies
Both LLM-inferred sources and sinks are necessary. Table 3 presents additional results when using either only the source or sink specification from an LLM in IRIS. For this experiment, we only use the results with GPT-4 for comparison. Each row present the number of detected vulnerabilities per CWE. We observe that omitting either source or sink specifications inferred by GPT-4 causes a drastic reduction in overall recall.
Performance gain of contextual analysis depends on LLM’s reasoning capability. As shown in Fig. 9, contextual analysis is highly necessary for the precision and F1 score improvements. However, only GPT-4, GPT-3.5, and Llama-3 70B see a positive impact after contextual analysis, while the smaller models see negative. The false positive reduction of contextual analysis is the most effective when the LLM possesses decent reasoning capability. Indeed, smaller models are more likely to respond with “vulnerable” than larger models.
CWE | 22 | 78 | 79 | 94 | Total |
22 | 1 | 4 | 0 | 27 () | |
28 | 3 | 5 | 0 | 36 () | |
10 | 1 | 9 | 4 | 24 () | |
31 | 3 | 13 | 8 | 55 |